A $25 shirt was marked up 20% then discounted by 20%. What is the new price of the shirt?
Mr.Cameron deposited $2,000 into a bank account, earning 3% simple interest for 5 years. How much money is in the account after that period of time?
Mr.Egelko is paid a 15% comission for a math game he sells. What is his comission on a game that sells for $25?
Ms.

Answer by checkley77(12844)

25*1.2=30
30*.8=24 PRICE OF THE SHIRT AFTER THE DISCOUNT.
-------------------------------------------------
I=(PRT)=2,000*.03*5=2,000*.15=300 IS THE INTEREST AFTER 5 YEARS.
THUS THE ACCOUNT IS WORTH 2,000+300=$2,300
------------------------------------------------
25*.15=$3.75 IS HIS COMMISSION ON EACH BOOK.